# PAT: Accelerating LLM Decoding via Prefix-Aware Attention with Resource Efficient Multi-Tile Kernel

### v2

@@ Abstract.
Large Language Models, LLM Inference, Prefix Aware Attention, GPU Kernel Scheduling
copyright: rightsretainedjournalyear: 2026conference: 31th ACM International Conference on Architectural Support for Programming Languages and Operating Systems; March 22–March 26, 2026; Pittsburgh, USAbooktitle: Proceedings of the 31th ACM International Conference on Architectural Support for Programming Languages and Operating Systems (ASPLOS ’26), March 32–March 26, 2026, Pittsburgh, USAdoi: 10.0000/0000000.0000000isbn: 000-0-0000-0000-0/00/00ccs: Computing methodologies Machine learningccs: Computer systems organization Cloud computing
@@ 1. Introduction
Tranformer-based Large Language Models (LLMs) have rapidly advanced across diverse domains, including conversational assistants (dam2024complete; dong2023towards), code generation (anthropic2025claudecode; hui2024qwencoder), retrieval-augmented generation (RAG) (lewis2020retrieval; yao2025cacheblend), and agent/tool workflows (microsoft2025autogen; yuan2024easytool). As deployment grows, latency and resource efficiency of online inference have become increasingly critical. While prior work has explored optimizations in memory management (kwon2023pagedattention; xu2025ellm), scheduling (zhong2024distserve; sun2024llumnix), quantization (lin2024awq; chee2023quip), and architecture (zhou2022mixture; liu2024deepseek), there remain new challenges and opportunities. We highlight two emerging trends below.
Longer contexts and outputs. Context length has scaled to millions of tokens (meta2025llama4), while techniques like Chain-of-Thought (CoT) (wei2022chain) demand longer outputs. As a result, decode attention (i.e., attention operations in the decoding stage) increasingly dominates end-to-end latency due to repeated loading of growing KV caches from global to on-chip memory, which poses a memory-bound challenge.
Shared prefixes across requests. System prompts, RAG documents, and agent templates introduce multi-level shared prefixes across requests, creating an optimization opportunity. Existing systems (zheng2024sglang; kwon2023pagedattention) implement prefix KV cache reuse, which reduces memory footprint by storing and reusing KV cache across requests. However, it can not reduce global memory accesses, which is the bottleneck for decode attention, thereby leading to higher attention latency.
Recent attention kernel implementations aim to reduce attention latency (shah2024flashattention3; ye2025flashinfer; ye2024chunkattention; pan2025fasttree; song2024tackling; zhu2024relayattention), but still face redundant KV cache loads from global memory and inefficient resource utilization. Query-centric kernels (dao2022flashattention; dao2023flashattention2; ye2025flashinfer; song2024tackling) follow a one-query-per-CTA strategy to map each query and corresponding KV cache into an independent Cooperative Thread Array (CTA; also known as a thread block) for execution on the GPU, causing redundant global memory loads for shared prefixes. KV-centric kernels (pan2025fasttree; zhu2024relayattention; yao2024deft; ye2024chunkattention) instead pack the KV cache of cross-query shared prefix into one CTA to reduce redundant memory accesses, but adopt a one-size-fits-all design that fixes the tile size of the GPU kernel (nvidia2025cudabestpractices) and applies padding to fill the tile, which wastes on-chip memory and limits CTA concurrency.
@@
To address these challenges, we present PAT, an attention kernel implementation for LLM decode stage with a pack–forward–merge execution paradigm. Specifically, in the pack stage5), PAT employs a prefix-aware pack scheduler and a lazy update mechanism, which efficiently pack decode batches into CTAs to reduce redundant global memory accesses with negligible overhead. It further designs a multi-tile kernel based on resource-efficiency analysis, so as to adapt to CTA dynamicity and avoid on-chip memory waste. In the forward stage6), PAT designs a multi-stream forward and long-KV split strategy, which parallelizes the multi-tile kernel across multiple CUDA streams and splits CTAs with excessively long KV lengths. These designs enable PAT to eliminate execution bubbles and achieve high global memory bandwidth utilization. In the merge stage7), PAT applies a lightweight kernel with online softmax (dao2022flashattention), which merges partial results across CTAs for each query.
@@
    We evaluate PAT on synthetic and real-world workloads, demonstrating that compared with state-of-the-art baselines, it reduces attention latency by 67.4% on average under the same decode batches (§8.3) and lowers TPOT by 17.0-93.1% under the same request rate (§8.4).
@@ 2.1. LLM Inference and Attention
Figure 1. Latency breakdown of Llama-3-8B and Qwen3-8B across context length on A100, vLLM v0.9.0, batch size 64:
@@ 2.2. Memory Bottleneck of Attention
Driven by the inference scaling law (wu2025inference), LLMs are trending toward longer contexts and longer outputs. For instance, Llama-4 supports up to 10 million input tokens (meta2025llama4), and Chain-of-Thought prompting (wei2022chain) has significantly increased output lengths for complex reasoning.
Although KV Caching avoids repeated computation, it has shifted the bottleneck from compute to memory access: each decode step must fetch a growing amount of cached KK/$V$ data from global memory to on-chip memory (zadouri2025hardware; behnam2025rocketkv). As input and output lengths grow, this access cost dominates inference latency. For models like Llama-3-8B and Qwen3-8B, decode attention can contribute up to 68% of the total latency, as shown in Figure 1.
@@ 2.3. GPU Execution Model
Figure 2. GPU architecture of typical NVIDIA GPUs.:
@@
Table 1. Memory hierarchy of A100-SXM4-80GB (nvidia2020a100; abdelkhalik2022demystifying).
Figure 3. (a) Two General Matrix-Vector multiplications (GEMV) in attention. (b) Tiling execution pipeline.:
When executing an attention kernel, the workload is divided into CTAs, scheduled across SMs. During this process, the required KV cache data is transferred from slow global memory up the hierarchy into fast on-chip shared memory, where it is finally loaded into registers for computation. State-of-the-art implementations (dao2022flashattention; dao2023flashattention2; ye2025flashinfer) are designed to exploit this memory hierarchy. For example, FlashAttention (dao2022flashattention; nvidia2025cudabestpractices) partitions K/V caches into small tiles along the sequence length (Figure 3a) and processes them in a pipelined fashion: while one tile is computed, the next is prefetched asynchronously (Figure 3b), reducing memory latency.
@@ 3. Motivation
In this section, we first characterize the shared-prefix pattern commonly observed in LLM workloads (§3.1) and identify two limitations in existing attention implementations under this pattern: redundant global memory accesses (§3.2) and low hardware utilization (§3.3). We then introduce the pack–forward–merge paradigm to address these problems (§3.4) and analysis the corresponding challenges (§3.5).
@@ 3.1. Shared Prefixes and Prefix Reuse
Shared prefixes are prevalent in modern LLM workloads (openai2025promptcaching; wang2020generalizing; jin2024dvd), with studies showing that 40–62% of KV cache prefixes are shared across requests (qin2025mooncake; wang2025kvcache). These prefixes often form multi-level hierarchies (zheng2024sglang; pan2025fasttree). For example, a base system prompt shared by all users, followed by task-specific prefixes shared within subgroups.
Systems like SGLang(zheng2024sglang) and vLLM(kwon2023pagedattention) support prefix reuse by mapping shared logical prefixes to a single physical copy in GPU global memory. This reduces KV cache memory usage and improves concurrency. However, it can not reduce global memory access, which remains the decode attention bottleneck. To understand why, we analyze two key inefficiencies in current attention kernels under shared-prefix scenarios: redundant memory access (§3.2) and underutilized hardware resources (§3.3) as follows.
@@ 3.2. Redundant Memory Access
A major inefficiency in current attention kernels stems from redundant memory accesses. To execute a decode batch, attention kernels use a packing strategy that groups queries and their KV caches into CTAs. In the decode batch with 4 queries shown in Figure 4a, existing kernels’ query-centric paradigm adopts a one-query-per-CTA packing strategy (shah2024flashattention3; ye2025flashinfer), where each query and its KV blocks are independently assigned to a CTA (Figure 4b). While simple to schedule, this strategy causes shared KV prefixes (e.g., KV-0, KV-1) to be repeatedly loaded from slow global memory into on-chip memory. Although L2 cache offers partial reuse, it is limited by size and bandwidth.
Figure 4. Comparison of packing strategies. (a) A decode batch of 4 queries with shared prefixes. (b) Query-centric packing causes redundant memory access. (c) KV-centric packing causes memory waste. (d) Memory-centric prefix-aware packing avoids redundancy and improves utilization.:
Figure 5. Limitations of existing attention kernels. (a) Average KV cache load from global memory per decode step on the toolagent and conversation traces, comparing FlashAttention (FA), PAT, and the theoretical optimum. (b) Two-dimensional resource inefficiencies.:
@@ 3.3. Two-Dimension Resource Inefficiencies
Furthermore, existing attention kernels exhibit inefficient hardware utilization. Both query-centric and KV-centric designs use a tiled pipeline (Figure 3) to overlap memory access and computation. In this paradigm, selecting appropriate tile sizes (i.e., query tile size $m$ and KV tile size $n$ from Figure 3b) is critical for resource efficiency (dao2022flashattention). However, existing kernels adopt a one-size-fits-all design, which specifies a single, hard-coded tile size for all CTAs (e.g., $m=64,n=32$ (shah2024flashattention3)).
As shown in Figure 4b-c, this static approach ignores the dynamic nature of LLM workloads, leading to significant resource inefficiencies along two dimensions [^1] While specific prior works (pan2025fasttree) design one more query tile sizes $m$, it is still insufficient for handling highly dynamic query numbers and KV lengths. in Figure 5(b): (1) Memory Waste ($\mathbb{I}_{mem}$): When fewer than $m$ queries share a KV prefix, CTAs must pad inputs, wasting shared memory and registers to store unused data. (2) Execution Bubble ($\mathbb{I}_{exe}$): Due to varying KV lengths across CTAs, fixed-size tiling causes imbalanced workloads, leaving SMs underutilized in the tail stages of execution.
@@ 3.4. Insight: Pack-forward-merge Paradigm
Following these principles, we propose a pack–forward–merge paradigm as follows: (1) Pack: group queries by shared prefix into CTAs to eliminate redundant KV loads; (2) Forward: execute CTAs using resource-optimized kernels that output partial results in tiles; (3) Merge: apply online softmax to combine partial results into final outputs.
@@ 4. Overview
To address these challenges, we design PAT, a memory-centric attention kernel implementation that follows the pack-forward-merge paradigm. It serves as a backend for the serving system vLLM (kwon2023pagedattention). In the pack stage, PAT adopts a profit-model-based heuristic packing strategy to aggregate queries that share KV into a CTA, so as to mitigate redundant global memory accesses (§5.1). It further designs multi-tile kernels and a runtime tile selector to choose an efficient kernel for each CTA (§5.2). In the forward stage, PAT adopts multi-stream execution and a long-KV split strategy to enable efficient kernel execution, which reduces execution bubbles (§6). Finally, in the merge stage, PAT uses a lightweight kernel based on online softmax to merge each query’s intermediate results across CTAs and produce the final output (§7).
@@ 5. Pack Scheduler
Figure 6. Workflow of the pack scheduler: (a) An input decode batch with 4 queries. (b) Tree structure block table. (c) Packing scheme 1 that splits leaf nodes with the parent node. (d) Packing scheme 2 that merges specific leaf node with the parent node.:
@@ 5.1. Pack Scheduler
Problem Formulation Since decode attention is memory-bound (§2.3 and §3.2), the pack stage aims to minimize global memory accesses for a given decode batch (Figure 6a). The input is a batch of queries plus a block table where each row lists block IDs for a query; a shared prefix appears as identical leading block IDs across rows. The output is a partition $\mathcal{P}=\{P_{1},P_{2},\dots\}$ of CTAs, where each $P_{i}$ packs queries sharing successive identical prefix blocks (queries may be split across CTAs). We seek $\mathcal{P}^{\ast}$ that minimizes memory accesses, comprised of loaded KV cache size and per-query intermediate reads/writes due to splits/merges. The search space grows exponentially with query count and prefix lengths  (pan2025fasttree), so an exact solver is impractical for online serving, and we turn to a heuristic pack scheduler described next.
@@
Intra-node profit. We first discuss the profit and overhead of packing the queries within a single non-leaf node $u$ (with attributes $l_{u}$ and $s_{u}>1$) into one CTA. First, compared with the one-query-per-CTA paradigm, packing $s_{u}$ queries with shared KV length of $l_{u}$ into a single CTA could reduce the KV cache loads from $s_{u}$ times to $1$ time. Therefore saves $(s-1)\,l_{u}\,d$ global memory accesses ($d$ is the head dimension). While packing the shared prefixes of all queries into a CTA produces $2\,s_{u}$ times (half comes from node $u$ and the other half from leaf nodes) per-query intermediate writes and reads. Therefore the overhead of memory access [^2] To ensure numerical accuracy, intermediate results are stored in FP32, so the overhead is multiplied by 2. is $2\times(2\,s_{u}\,d+2\,s_{u}\,d)\,=\,8\,s_{u}\,d$. Then, the profit–overhead ratio for packing a non-leaf node to a CTA is:
@@
In practice, the length of shared KV $l_{u}\geq 16$ since sharing is performed at the granularity of KV blocks (kwon2023pagedattention; shah2024flashattention3), whose sizes are typically larger than 16. Therefore, packing a node into a CTA yields a positive profit.
@@
Lazy Update. Although the pack scheduler is linear, its overhead can grow with batch size and number of KV blocks. We mitigate this with a lazy-update strategy that (1) reuses a scheduling result across continuous-batching iterations until the block table changes (e.g., request arrivals/departures or new KV block assignments) and (2) moves the scheduler into the serving system and runs it asynchronously once the block table is available. These reduce scheduler invocations from once per transformer layer to once several continuous-batching iterations and overlap scheduling with pre-attention task (e.g., LayerNorm, QKV projection), thereby substantially reducing the exposed scheduling latency (see §8.6) without affecting model accuracy.

### v3

@@ Abstract.
Large Language Models; LLM Inference; Prefix Aware Attention; GPU Kernel Scheduling
journalyear: 2026copyright: ccconference: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2; March 21–26, 2026; Pittsburgh, PA, USA.booktitle: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2 (ASPLOS ’26), March 21–26, 2026, Pittsburgh, PA, USAisbn: 979-8-4007-2359-9/2026/03doi: 10.1145/3779212.3790200ccs: Computing methodologies Machine learningccs: Computer systems organization Cloud computing
@@ 1. Introduction
Transformer-based Large Language Models (LLMs) have rapidly advanced across diverse domains, including conversational assistants (Dam et al., 2024; Dong et al., 2023), code generation (Anthropic, 2025; Hui et al., 2024), retrieval-augmented generation (RAG) (Lewis et al., 2020; Yao et al., 2025), and agent/tool workflows (Microsoft, 2025; Yuan et al., 2024). As deployment grows, latency and resource efficiency of online inference have become increasingly critical. While prior work has explored optimizations in memory management (Kwon et al., 2023; Xu et al., 2025), scheduling (Zhong et al., 2024; Sun et al., 2024), quantization (Lin et al., 2024b; Chee et al., 2023), and architecture (Zhou et al., 2022; Liu et al., 2024), there remain new challenges and opportunities. We highlight two emerging trends below.
Longer contexts and outputs. Context length has scaled to millions of tokens (AI, 2025), while techniques like Chain-of-Thought (CoT) (Wei et al., 2022) demand longer outputs. As a result, decode attention (i.e., attention operations in the decoding stage) increasingly dominates end-to-end latency due to repeated loading of growing KV caches from global to on-chip memory, which poses a memory-bound challenge.
Shared prefixes across requests. System prompts, RAG documents, and agent templates introduce multi-level shared prefixes across requests, creating an optimization opportunity. Existing systems (Zheng et al., 2024b; Kwon et al., 2023) implement prefix KV cache reuse, which reduces memory footprint by storing and reusing KV cache across requests. However, it can not reduce global memory accesses, which is the bottleneck for decode attention, thereby leading to higher attention latency.
Recent attention kernel implementations aim to reduce attention latency (Shah et al., 2024; Ye et al., 2025, 2024b; Pan et al., 2025; Song et al., 2024; Zhu et al., 2024), but still face redundant KV cache loads from global memory and inefficient resource utilization. Query-centric kernels (Dao et al., 2022; Dao, 2023; Ye et al., 2025; Song et al., 2024) follow a one-query-per-CTA strategy to map each query and corresponding KV cache into an independent Cooperative Thread Array (CTA; also known as a thread block) for execution on the GPU, causing redundant global memory loads for shared prefixes. KV-centric kernels (Pan et al., 2025; Zhu et al., 2024; Yao et al., 2024; Ye et al., 2024b) instead pack the KV cache of cross-query shared prefix into one CTA to reduce redundant memory accesses, but adopt a one-size-fits-all design that fixes the tile size of the GPU kernel (NVIDIA, 2025a) and applies padding to fill the tile, which wastes on-chip memory and limits CTA concurrency.
@@
To address these challenges, we present PAT, an attention kernel implementation for LLM decode stage with a pack–forward–merge execution paradigm. Specifically, in the pack stage5), PAT employs a prefix-aware pack scheduler and a lazy update mechanism, which efficiently pack decode batches into CTAs to reduce redundant global memory accesses with negligible overhead. It further designs a multi-tile kernel based on resource-efficiency analysis, so as to adapt to CTA dynamicity and avoid on-chip memory waste. In the forward stage6), PAT designs a multi-stream forward and long-KV split strategy, which parallelizes the multi-tile kernel across multiple CUDA streams and splits CTAs with excessively long KV lengths. These designs enable PAT to eliminate execution bubbles and achieve high global memory bandwidth utilization. In the merge stage7), PAT applies a lightweight kernel with online softmax (Dao et al., 2022), which merges partial results across CTAs for each query.
@@
    We evaluate PAT on synthetic and real-world workloads, demonstrating that compared with state-of-the-art baselines, it reduces attention latency by 53.5% on average under the same decode batches (§8.3) and lowers TPOT by 17.0-93.1% under the same request rate (§8.4).
@@ 2.1. LLM Inference and Attention
Figure 1. Latency breakdown of Llama-3-8B and Qwen3-8B across context length on A100, vLLM v0.9.0, batch size 64
@@ 2.2. Memory Bottleneck of Attention
Driven by the inference scaling law (Wu et al., 2025), LLMs are trending toward longer contexts and longer outputs. For instance, Llama-4 supports up to 10 million input tokens (AI, 2025), and Chain-of-Thought prompting (Wei et al., 2022) has significantly increased output lengths for complex reasoning.
Although KV Caching avoids repeated computation, it has shifted the bottleneck from compute to memory access: each decode step must fetch a growing amount of KV cache from global memory to on-chip memory (Zadouri et al., 2025; Behnam et al., 2025). As input and output lengths grow, this access cost dominates inference latency. As in Figure 1, decode attention can contribute up to 53% of the total latency for Llama-3-8B and Qwen3-8B.
@@ 2.3. GPU Execution Model
Figure 2. GPU architecture of typical NVIDIA GPUs.
@@
Table 1. Memory hierarchy of A100-SXM4-80GB (NVIDIA, 2020; Abdelkhalik et al., 2022).
Figure 3. (a) Two General Matrix-Vector multiplications (GEMV) in attention. (b) Tiling execution pipeline.
When executing an attention kernel, the workload is divided into CTAs, scheduled across SMs. During this process, the required KV cache data is transferred from slow global memory up the hierarchy into fast on-chip shared memory, where it is finally loaded into registers for computation. State-of-the-art implementations (Dao et al., 2022; Dao, 2023; Ye et al., 2025) are designed to exploit this memory hierarchy. For example, FlashAttention (Dao et al., 2022; NVIDIA, 2025a) partitions K/V caches into small tiles along the sequence length (Figure 3a) and processes them in a pipelined fashion: while one tile is computed, the next is prefetched asynchronously (Figure 3b), reducing memory latency.
@@ 3. Motivation
In this section, we first characterize the shared-prefix pattern commonly observed in LLM workloads (§3.1) and identify two limitations in existing attention implementations under this pattern: redundant global memory accesses (§3.2) and low hardware utilization (§3.3). We then introduce the pack–forward–merge paradigm to address these problems (§3.4) and analyze the corresponding challenges (§3.5).
@@ 3.1. Shared Prefixes and Prefix Reuse
Figure 4. Prefix ratio of four traces: ToolAgent: tool and agent interaction workload (Qin et al., 2025); Conversation: online conversation workload (Qin et al., 2025); Qwen-A: online API service (Wang et al., 2025); Qwen-B: task automation with API calling workload (Wang et al., 2025).
Shared prefixes between requests are common and hierarchical in modern LLM workloads (OpenAI, 2025; Wang et al., 2020; Jin et al., 2024). As shown in Figure 4, our analysis of four real-world traces (Qin et al., 2025; Wang et al., 2025) finds a prefix ratio of $51.9-75.0\%$, meaning that more than half of KV cache tokens come from prefixes reused across requests. Under continuous batching, this cross-request sharing leads to multi-level intra-batch prefixes. On the Conversation and ToolAgent traces, using the setup in §8.2, intra-batch shared prefixes cover $2.8-82.6\%$ of KV caches, and each batch ends up with 2.72 distinct shared prefixes on average. These significant and hierarchical shared prefixes present unique opportunities and challenges for LLM inference optimization.
Systems like SGLang(Zheng et al., 2024b) and vLLM(Kwon et al., 2023) adopt prefix reuse by mapping shared logical prefixes to a single physical copy in global memory. This approach reuses KV caches across requests to reduce memory usage. However, it can not leverage intra-batch shared prefixes to reduce global memory accesses, which remain the bottleneck. To understand why, we analyze two key inefficiencies in existing approaches under shared-prefix scenarios: redundant memory access (§3.2) and underutilized hardware resources (§3.3).
@@ 3.2. Redundant Memory Access
A major inefficiency in current attention kernels stems from redundant memory accesses. To execute a decode batch, attention kernels use a packing strategy that groups queries and their KV caches into CTAs. In the decode batch with 4 queries (Figure 5a), existing kernels’ query-centric paradigm adopts a one-query-per-CTA packing strategy (Shah et al., 2024; Ye et al., 2025), where each query and its KV are independently assigned to a CTA (Figure 5b). While simple to schedule, this strategy causes shared KV prefixes (e.g., KV-0, KV-1) to be repeatedly loaded from slow global memory into on-chip memory. Although L2 cache offers partial reuse, it is limited by size and bandwidth.
Figure 5. Comparison of packing strategies. (a) A decode batch of 4 queries with shared prefixes. (b) Query-centric packing causes redundant memory access. (c) KV-centric packing causes memory waste. (d) Memory-centric prefix-aware packing avoids redundancy and improves utilization.
Figure 6. Limitations of existing attention kernels. (a) Average KV cache load from global memory per decode step on the toolagent and conversation traces, comparing FlashAttention (FA), PAT, and the theoretical optimum. (b) Two-dimensional resource inefficiencies.
@@ 3.3. Two-Dimension Resource Inefficiencies
Furthermore, existing attention kernels exhibit inefficient hardware utilization. Both query-centric and KV-centric designs use a tiled pipeline (Figure 3) to overlap memory access and computation. In this paradigm, selecting appropriate tile sizes (i.e., query tile size $m$ and KV tile size $n$ from Figure 3b) is critical for resource efficiency (Dao et al., 2022). However, existing kernels adopt a one-size-fits-all design, which specifies a single, hard-coded tile size for all CTAs (e.g., $m=64,n=32$ (Shah et al., 2024)).
As shown in Figure 5b-c, this static approach ignores the dynamic nature of LLM workloads, leading to significant resource inefficiencies along two dimensions [^1] While specific prior works (Pan et al., 2025) design one more query tile sizes $m$, it is still insufficient for handling highly dynamic query numbers and KV lengths. in Figure 6(b): (1) Memory Waste ($\mathbb{I}_{mem}$): When fewer than $m$ queries share a KV prefix, CTAs must pad inputs, wasting shared memory and registers to store unused data. (2) Execution Bubble ($\mathbb{I}_{exe}$): Due to varying KV lengths across CTAs, fixed-size tiling causes imbalanced workloads, leaving SMs underutilized in the tail stages of execution.
@@ 3.4. Insight: Pack-forward-merge Paradigm
Following these principles, we propose pack–forward–merge paradigm as follows: (1) Pack: group queries by shared prefix into CTAs to eliminate redundant KV loads; (2) Forward: execute CTAs using resource-optimized kernels that output partial results in tiles; (3) Merge: apply online softmax to combine partial results into final outputs.
@@ 4. Overview
To address these challenges, we design PAT, a memory-centric attention kernel implementation that follows the pack-forward-merge paradigm. It serves as a backend for the serving system vLLM (Kwon et al., 2023). In the pack stage, PAT adopts a profit-model-based heuristic packing strategy to aggregate queries that share KV into a CTA, so as to mitigate redundant global memory accesses (§5.1). It further designs multi-tile kernels and a runtime tile selector to choose an efficient kernel for each CTA (§5.2). In the forward stage, PAT adopts multi-stream execution and a long-KV split strategy to enable efficient kernel execution, which reduces execution bubbles (§6). Finally, in the merge stage, PAT uses a lightweight kernel based on online softmax to merge each query’s intermediate results across CTAs and produce the final output (§7).
@@ 5. Pack Scheduler
Figure 7. Workflow of the pack scheduler: (a) An input decode batch with 4 queries. (b) Tree structure block table. (c) Packing scheme 1 that splits leaf nodes with the parent node. (d) Packing scheme 2 that merges specific leaf node with the parent node.
@@ 5.1. Pack Scheduler
Problem Formulation Since decode attention is memory-bound (§2.3 and §3.2), the pack stage aims to minimize global memory accesses for a given decode batch (Figure 7a). The input is a batch of queries plus a block table where each row lists block IDs for a query; a shared prefix appears as identical leading block IDs across rows. The output is a partition $\mathcal{P}=\{P_{1},P_{2},\dots\}$ of CTAs, where each $P_{i}$ packs queries sharing successive identical prefix blocks (queries may be split across CTAs). We seek $\mathcal{P}^{\ast}$ that minimizes memory accesses, comprised of loaded KV cache size and per-query intermediate reads/writes due to splits/merges. The search space grows exponentially with query count and prefix lengths  (Pan et al., 2025), so an exact solver is impractical for online serving, and we turn to a heuristic pack scheduler described next.
@@
Intra-node profit. We first discuss the profit and overhead of packing the queries within a single non-leaf node $u$ (with attributes $l_{u}$ and $s_{u}>1$) into one CTA. First, compared with the one-query-per-CTA paradigm, packing $s_{u}$ queries with shared KV length of $l_{u}$ into a single CTA could reduce the KV cache loads from $s_{u}$ times to $1$ time. Therefore, it saves $(s-1)\,l_{u}\,d$ global memory accesses ($d$ is the head dimension). However, packing the shared prefixes of all queries into a CTA produces $2\,s_{u}$ times (half comes from node $u$ and the other half from leaf nodes) per-query intermediate writes and reads. Therefore the overhead of memory access [^2] To ensure numerical accuracy, intermediate results are stored in FP32, so the overhead is multiplied by 2. is $2\times(2\,s_{u}\,d+2\,s_{u}\,d)\,=\,8\,s_{u}\,d$. Then, the profit–overhead ratio for packing a non-leaf node to a CTA is:
@@
In practice, the length of shared KV $l_{u}\geq 16$ since sharing is performed at the granularity of KV blocks (Kwon et al., 2023; Shah et al., 2024), whose sizes are typically larger than 16. Therefore, packing a node into a CTA yields a positive profit.
@@
Lazy Update. Although the pack scheduler is linear, its overhead can grow with batch size and number of KV blocks. We mitigate this with a lazy-update strategy that (1) reuses a scheduling result across continuous-batching iterations until the block table changes (e.g., request arrivals/departures or new KV block assignments) and (2) moves the scheduler into the serving system and runs it asynchronously once the block table is available. These reduce scheduler invocations from once per transformer layer to once several continuous-batching iterations and overlap scheduling with pre-attention task (e.g., metadata preparation, QKV projection), thereby substantially reducing the exposed scheduling latency (see §8.7) without affecting model accuracy.